# Also at ]Alpine Quantum Technologies GmbH, Innsbruck, Austria

### v1

@@ III Trapped-ion noisy Rabi oscillations: tomography and randomized benchmarking
Up to this point we have gauged the validity of our analytical predictions against numerical simulations with an ad hoc noise model, namely the extended phase diffusion model with an Ornstein-Uhlenbeck process $\delta\tilde{\omega}(t)$. We are now at the point where we can gauge the utility of our approach against more realistic noise models that include (structured) technical contributions. When deriving the analytical expressions for the process matrix and the effective Kraus operators, we have only assumed that the noise is wide-sense (or independent increment) stationary. As we have discussed, in this scenario we can express the dynamical map in terms of the filtered integrals in Eq. (18), so it can describe any scenario in which the predominant noise of the laser can be characterized using a power spectral density [94].
@@ III.1 Experimental setup
The trapped-ion QIP used in our experiments features a macroscopic Paul trap within which linear strings of ${}^{40}$${}^{+}$ ions can be confined. The overall experimental system that was used to generate the majority of data in this study is described in [124]. Optical qubits are encoded in the ground state |$\ket{\textrm{4\,S}_{1/2},m_{J}=-1/2}$ and excited metastable state |$\ket{\textrm{3\,D}_{5/2},m_{J}=-1/2}$, which are connected via an electric quadrupole qubit transition near $729\ {\rm nm}$. The electric quadrupole transition is driven by a frequency-stabilized Titanium:Sapphire laser whose frequency fluctuations lead to the dephasing noise process $\delta\tilde{\omega}(t)$ in Eq. (8) that has been studied above. Projective measurements in the computational basis are performed by state-selective fluorescence. When qubits in the electronic |$\ket{\textrm{4\,S}_{1/2},m_{J}=-1/2}$ state are illuminated by a readout beam near $397\,{\rm nm}$ the states will be rapidly driven on an electric dipole transition, scattering many photons which can be collected on a camera. If the qubit is however in the |$\ket{\textrm{3\,D}_{5/2},m_{J}=-1/2}$ they will not participate in this transition, such that the qubit state can be inferred by the presence (absence) of substantial fluorescence signal. We use a simple threshold-based discriminator between the two qubit states which leads to measurement infidelity better than $10^{-3}$ [124].
@@
Figure 6: Experimental laser frequency power spectral density with (blue) and without (orange) measurement artifacts. The broad peak at $\omega\approx 400\cdot 10^{3}\,{\rm rad/s}$ is a servo bump not present at the qubit location.
@@ III.2 Noisy tomography of the dynamical error map and validation of the analytical error rate estimates
We extract the experimental gate error from the measurement data by applying the QPT framework discussed in Sec.  II.3 to the set of experimentally obtained relative frequencies. The presence of further sources of noise in the data necessitates use of the log-likelihood estimator in Eq. (42) instead of linear inversion to obtain only physically admissible channels. However, statistical treatment presented above would lead to poor estimates for mean gate error and variance given the limited amount of shots per time we can obtain in experiments. It would not be sensible to further subdivide the data to make statistical predictions of the confidence regions. Therefore, we have followed the alternative method described in [122, 123]. This method becomes unwieldy for large amounts of data, but its great advantage is direct access to the statistics of the distribution. Thus, within our regime of sparse data it can more accurately and efficiently predict confidence intervals.
To allow for an efficient Monte Carlo sampling of the distribution of the parameters, one must consider directly a rescaled multinomial likelihood function rather than the previous negative log-likelihood (42), that is
@@
which is proportional to the confidence that the underlying quantum channels for the snapshots are described by a given process matrix $\chi(t_{i})$ given the experimental data set $\mathcal{D}=\{n_{s,\mu,\nu}(t_{i})\}$. Here, $n_{s,\mu,\nu}(t_{i})\in\{0,1\}$ are the specific binary outcomes for each POVM measurement and time-evolved initial state, taking into account the mutually-exclusive cases. Note that using $N_{s,i,\mu}=\sum_{\nu}n_{s,\mu,\nu}(t_{i})$, and taking the rescaled negative logarithm, connects to the previous cost function (42). This likelihood function peaks at the experimental relative frequency as before, but in contrast to the log-likelihood estimator, the dispersion of the distribution will narrow with the amount of available data making for a more precise prediction of the confidence regions.
Given a likelihood function, it is not easy to obtain the underlying probability distribution, as the scaling factor is generally unknown. However, we can reconstruct this distribution using an a Metropolis-Hasting acceptance-rejection Markov chain Monte Carlo method [125]. For each time step, this algorithm starts from an initial assumption of the process matrix $\chi(t_{i})$, and, given a proposed new value of the matrix, it either accepts or rejects it with a certain update probability $\alpha$ set by the ratio of the above likelihood function (46) evaluated at both the current and proposed values of the process error matrix. Note that if we use asymmetrical probability distributions to propose the new process error matrix from the current one, the above update probability needs to be corrected by the ratio of the corresponding probability distributions to not introduce a bias in the Monte Carlo updates. We impose the physical constraints of the process error matrix (42) in a different way to the maximum likelihood method described above to perform the Metropolis-Hastings algorithm efficiently in the context of QPT. In particular, we impose its positivity by resorting to the Cholesky decomposition, i.e., decomposing $\chi(t_{i})={L}_{\chi}(t_{i}){L}_{\chi}^{\dagger}(t_{i})$, where ${L}_{\chi}(t_{i})$ is a lower triangular matrix, and demanding that all of its diagonal entries are positive. Taking into account the block-diagonal structure of the microscopic process error matrix (see the discussion around Eq. (99) in Appendix D), we can reduce the number of entries of the lower triangular matrix as
@@
Hence, the reconstruction of the estimator distribution requires us to sample each entry of ${L}_{\chi}(t_{i})$ under the above constraints for each iteration. We use a truncated Gaussian distribution in the interval $\ell_{\alpha\beta}(t_{i})\in[0,1]$ for the diagonal elements and $\ell_{\alpha\beta}(t_{i})\in(-1,1)$ for the non-diagonal ones. Furthermore, to be sure that the trace-preserving condition holds for all the proposed matrix elements, we need to normalize the proposed values using Eq. (48). To reconstruct the distribution of each element, we have used a Metropolis-Hastings chain of $M_{\rm MC}=10^{5}$ steps, which allows for sufficient smoothness in the reconstructed histograms that approximate the corresponding probability distribution of the estimated process matrix. Once we have these histograms, we can compute the corresponding gate error (45) to then calculate its distribution amenable to discrete statistics. In particular, we can compute the maximum of the distribution by computing the mode and the corresponding confidence region by computing the desired quantiles of the histogram.
Figure 7: Reconstruction of the gate error from experimental data and how it compares with the analytical predictions. In a we depict the 95% confidence region of the gate error (shaded area) for the estimated snapshots using QPT, as well as the maximum-likelihood estimate (purple markers, line as guide to the eye). The experimental data is subject to SPAM errors and shot noise associated with the $N_{i}=1200$ measurements per time. We depict the analytical error prediction of the gate error (45) (red) due to solely the filtered dephasing noise PSD of Fig. 6. In b we show the randomized benchmarking estimates for the gate error (purple) against our non-Markovian estimation (red) and the standard depolarizing model (green) for a $\pi$-pulse as a function of the Rabi frequency of the system. Error bars of the data here are below the marker size.
@@
In the Clifford gate decomposition used [127] there are an average of $\approx 2.2$ pulses which we then may use to find a proxy of average gate fidelity on the pulse level. Note that this fidelity estimate is somewhat different to the fidelity obtained from the dynamical error map. Firstly, the randomized benchmarking decay is not solely due to phase noise, which means the estimated error rate is likely higher than those considering only dephasing. Secondly, the dynamical error map estimate uses a fixed Pauli axis in the evolution but averages over input states instead. An additional difference is that randomized benchmarking is known to provide inaccurate fidelity estimates in the presence of correlated errors [128], whereas the method presented here is robust to this. Lastly, $\epsilon_{\text{RB}}/2.2$ is not a theoretically rigorous estimate of an average gate fidelity on the pulse level, but we include it as it is very commonly used in the community. We then compare the theoretical prediction for a $\pi$-pulse and the randomized benchmarking estimate for a range of Rabi frequencies that can be achieved experimentally. The benchmarking estimates at the different Rabi frequencies are lower-bounded by our analytical estimates which can be seen in Fig. 7b. This is expected given our treatment only consider phase noise when the experiment will also experience other sources of noise. They are notably below the standard infidelity estimates commonly used in the literature.
@@ IV Conclusions and outlook
In this work we have provided a microscopically motivated tomography of the dynamical error map of single-qubit gates subject to colored dephasing noise. We have provided an efficient parametrization of the dynamical error map that depends only on various filtered integrals of the underlying noise power spectral density. By working in a dressed-state instantaneous frame, we have shown that the cumulant expansion in the time-convolutionless master equation approach can be truncated at second order and yet capture the non-Markovian dynamics of the stochastic quantum system in a compact completely-positive and trace-preserving map whose accuracy substantially outperforms a more standard depolarizing map. Our derived dynamical error map, valid at arbitrary times, is in general not a Clifford-type channel, and provides predictive power based solely on experimentally accessible power spectral densities. Moreover, it allows to qualitatively and quantitatively identify non-Markovian contributions to the noise, all of which are condensed in various simple filtered integrals that, to the best of our knowledge, have not been previously discussed in the literature. We have also provided a compact Pauli-twirled approximation of the channel, which maintains the non-Markovian effect during each driving within a single gate, but can be used in large-scale simulations of noisy Clifford circuits.
@@
We have gauged our analytical approach extensively against direct numerical simulations of the noisy dynamics as well as against map reconstructions based on various approximate models. Comparing the average channel infidelity of single-qubit gates subject to an extended phase diffusion model of the laser, we have found that our non-Clifford and non-Markovian dynamical map outperforms the literature-standard depolarizing noise model. We further see that coherent over-rotations present in the noise can be explained when including non-Markovian contributions to the noise that have previously not been described. Given the broad assumptions of our treatment we have also used this model’s predictions to compare against experimental data obtained via quantum process tomography and randomized benchmarking. For short times, that is the regime of few Rabi flops in which QIPs typically operate, we see that non-Markovian contributions arise, although they are small in comparison to other contributions to the error. However, we point out that it is expected that non-Markovian contributions will become important in longer sequences of gates as are encountered in algorithms or QCVV methods that employ signal amplification. Capturing these contributions will thus be important in understanding the cumulative effect of noise correlations in QIPs for QCVV and applications.
@@
In our trapped-ion experiments, we find that our reconstructed noise model provides a good lower bound for established infidelity estimators such as randomized benchmarking. In particular, we find that our model outperforms previous gate infidelity metrics based on filtered power spectral densities, and captures the main contributions of experimental infidelity even if it only considers phase noise on the drive. We note that the resource overhead of quantum process tomography reconstructions and estimation of gate infidelity is substantial for trapped ions, such that they were entirely dominated by projection noise in our study. However, in experimental platforms where experimental shots are more readily accumulated, such as in superconducting qubits, this should remain a viable approach. Our predictions are compared against (standard) randomized benchmarking, which can be inaccurate under several circumstances that are experimentally relevant. In the future, other techniques such as cycle benchmarking [129] could be employed for a more stringent test.
@@ Acknowledgements.
A.B. warmly thanks P. Zoller for sharing the lecture notes in [82], for interesting discussions regarding the fast-fluctuation expansion in multiplicative SDEs including the connection to relevant previous works for the dressed-state formalism in SDEs [89], as well as noting the existence of exact solutions for the driven qubit under a Ornstein-Uhlenbeck frequency noise. A.S. would like to thank M. Meth for his assistance with operating the machine to take randomized benchmarking and QPT data. The project leading to this publication has received funding from the US Army Research Office through Grant No. W911NF-21-1-0007. A.B acknowledges support from PID2021-127726NB- I00 (MCIU/AEI/FEDER, UE), from the Grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033, from the CSIC Research Platform on Quantum Technologies PTI-001, and from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101114305 (“MILLENION-SGA1” EU Project). R.F., A.S., M.G.B., Ch.D.M., and T.M. also acknowledge funding by the Austrian Science Fund (FWF), through the SFB BeyondC (FWF Grant-DOI 10.55776/F71), and the IQI GmbH. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.

### v3

@@ III Trapped-ion noisy Rabi oscillations: tomography and randomized benchmarking
Up to this point we have gauged the validity of our analytical predictions against numerical simulations with an ad hoc noise model, namely the extended phase diffusion model with an Ornstein-Uhlenbeck process $\delta\tilde{\omega}(t)$. We are now at the point where we can gauge the utility of our approach against more realistic noise models that include (structured) technical contributions. When deriving the analytical expressions for the process matrix and the effective Kraus operators, we have only assumed that the noise is wide-sense (or independent increment) stationary, and that there is no cross-correlation between amplitude and phase noise. As we have discussed, in this scenario we can express the dynamical map in terms of the filtered integrals in Eq. (18), so it can describe any scenario in which the predominant noise of the laser can be characterized using a power spectral density [94].
@@ III.1 Experimental setup
[…]
@@
Figure 7: Experimental power spectral densities for the laser. a Frequency PSD with (blue) and without (orange) measurement artifacts. The broad peak at $\omega\approx 400\cdot 10^{3}\,{\rm rad/s}$ is a servo bump not present at the qubit location. b Rabi frequency PSD obtained from intensity fluctuations using a photovoltage detector.
@@
The relative intensity noise PSDs were recorded using a transimpedance-amplified photodetector positioned after a beam pickup close to the ion trap. The signal’s spectral density is measured by a audio analyzer (spectrum analyzer for low frequencies). The voltage drop (or photocurrent) measured by the square-law detector is proportional to the optical power. The ratio of photovoltage to average voltage is therefore identical to the ratio of optical intensity fluctuations to optical mean intensity $\delta\tilde{I}_{L}(t)/I_{L,0}$. We note that relative intensity noise is frequently defined as the ratio of electrical powers, not optical powers. However, here optical powers and not their square are required.
@@ III.2 Noisy tomography of the dynamical error map and validation of the analytical error rate estimates
We extract the experimental gate error from the measurement data by applying the QPT framework discussed in Sec.  II.4 to the set of experimentally obtained relative frequencies. The presence of further sources of noise in the data necessitates use of the log-likelihood estimator in Eq. (53) instead of linear inversion to obtain only physically admissible channels. However, statistical treatment presented above would lead to poor estimates for mean gate error and variance given the limited amount of shots per time we can obtain in experiments. It would not be sensible to further subdivide the data to make statistical predictions of the confidence regions. Therefore, we have followed the alternative method described in [123, 124]. This method becomes unwieldy for large amounts of data, but its great advantage is direct access to the statistics of the distribution. Thus, within our regime of sparse data it can more accurately and efficiently predict confidence intervals.
To allow for an efficient Monte Carlo sampling of the distribution of the parameters, one must consider directly a rescaled multinomial likelihood function rather than the previous negative log-likelihood in Eq. (53), that is
@@
which is proportional to the confidence that the underlying quantum channels for the snapshots are described by a given process matrix $\chi(t_{i})$ given the experimental data set $\mathcal{D}=\{n_{s,\mu,\nu}(t_{i})\}$. Here, $n_{s,\mu,\nu}(t_{i})\in\{0,1\}$ are the specific binary outcomes for each POVM measurement and time-evolved initial state, taking into account the mutually-exclusive cases. Note that using $N_{s,i,\mu}=\sum_{\nu}n_{s,\mu,\nu}(t_{i})$, and taking the rescaled negative logarithm, connects to the previous cost function in Eq. (53). This likelihood function peaks at the experimental relative frequency as before, but in contrast to the log-likelihood estimator, the dispersion of the distribution will narrow with the amount of available data making for a more precise prediction of the confidence regions.
Given a likelihood function, it is not easy to obtain the underlying probability distribution, as the scaling factor is generally unknown. However, we can reconstruct this distribution using an a Metropolis-Hasting acceptance-rejection Markov chain Monte Carlo method [126]. For each time step, this algorithm starts from an initial assumption of the process matrix $\chi(t_{i})$, and, given a proposed new value of the matrix, it either accepts or rejects it with a certain update probability $\alpha$ set by the ratio of the above likelihood function in Eq. (59) evaluated at both the current and proposed values of the process error matrix. Note that if we use asymmetrical probability distributions to propose the new process error matrix from the current one, the above update probability needs to be corrected by the ratio of the corresponding probability distributions to not introduce a bias in the Monte Carlo updates. We impose the physical constraints of the process error matrix in Eq. (53) in a different way to the maximum likelihood method described above to perform the Metropolis-Hastings algorithm efficiently in the context of QPT. In particular, we impose its positivity by resorting to the Cholesky decomposition, i.e., decomposing $\chi(t_{i})={L}_{\chi}(t_{i}){L}_{\chi}^{\dagger}(t_{i})$, where ${L}_{\chi}(t_{i})$ is a lower triangular matrix, and demanding that all of its diagonal entries are positive. Taking into account the block-diagonal structure of the microscopic process error matrix (see the discussion around Eq. (115) in Appendix D), we can reduce the number of entries of the lower triangular matrix as
@@
Hence, the reconstruction of the estimator distribution requires us to sample each entry of ${L}_{\chi}(t_{i})$ under the above constraints for each iteration. We use a truncated Gaussian distribution in the interval $\ell_{\alpha\beta}(t_{i})\in[0,1]$ for the diagonal elements and $\ell_{\alpha\beta}(t_{i})\in(-1,1)$ for the non-diagonal ones. Furthermore, to be sure that the trace-preserving condition holds for all the proposed matrix elements, we need to normalize the proposed values using Eq. (61). To reconstruct the distribution of each element, we have used a Metropolis-Hastings chain of $M_{\rm MC}=10^{5}$ steps, which allows for sufficient smoothness in the reconstructed histograms that approximate the corresponding probability distribution of the estimated process matrix. Once we have these histograms, we can compute the corresponding gate error in Eq. (56) to then calculate its distribution amenable to discrete statistics. In particular, we can compute the maximum of the distribution by computing the mode and the corresponding confidence region by computing the desired quantiles of the histogram.
Figure 8: Reconstruction of the gate error from experimental data and how it compares with the analytical predictions. In a we depict the 95% confidence region of the gate error (shaded area) for the estimated snapshots using QPT, as well as the maximum-likelihood estimate (purple markers, line as guide to the eye). The experimental data is subject to SPAM errors and shot noise associated with the $N_{i}=1200$ measurements per time. We depict the analytical error prediction of the gate error in Eq. (56) due to solely the filtered dephasing noise PSD of Fig. 7 (red), and with the amplitude fluctuations contribution (black). In b we show the randomized benchmarking estimates for the gate error (purple) against our non-Markovian phase-only estimation (red), the model including amplitude noise (black), and the standard depolarizing model (green) for a $\pi$-pulse as a function of the Rabi frequency of the system. Error bars of the data here are below the marker size.
@@
In the Clifford gate decomposition used [128] there are an average of $\approx 2.2$ pulses which we then may use to find a proxy of average gate fidelity on the pulse level. Note that this fidelity estimate is somewhat different to the fidelity obtained from the dynamical error map. Firstly, the randomized benchmarking decay is not solely due to the zero-mean phase and amplitude noise without cross-correlation, i.e. we expect some coherent (miscalibration) errors and potentially cross-correlation. The estimated error rate is therefore likely higher. Secondly, the dynamical error map estimate uses a fixed Pauli axis in the evolution but averages over input states instead. An additional difference is that randomized benchmarking is known to provide inaccurate fidelity estimates in the presence of (temporally) correlated errors [129], whereas the method presented here is robust to this. Lastly, $\epsilon_{\text{RB}}/2.2$ is not a theoretically rigorous estimate of an average gate fidelity on the pulse level, but we include it as it is very commonly used in the community. We then compare the theoretical prediction for a $\pi$-pulse and the randomized benchmarking estimate for a range of Rabi frequencies that can be achieved experimentally. The benchmarking estimates at the different Rabi frequencies are lower-bounded by our analytical estimates which can be seen in Fig. 8b. They are notably below the standard infidelity estimates commonly used in the literature.
@@ IV Conclusions and outlook
In this work we have provided a microscopically motivated tomography of the dynamical error map of single-qubit gates subject to colored multi-axis noise. We have provided an efficient parametrization of the dynamical error map that depends only on various filtered integrals of the underlying noise power spectral densities. By working in a dressed-state instantaneous frame, we have shown that the cumulant expansion in the time-convolutionless master equation approach can be truncated at second order and yet capture the non-Markovian dynamics of the stochastic quantum system in a compact completely-positive and trace-preserving map whose accuracy substantially outperforms a more standard depolarizing map. Our derived dynamical error map, valid at arbitrary times, is in general not a Clifford-type channel, and provides predictive power based solely on experimentally accessible power spectral densities. Moreover, it allows to qualitatively and quantitatively identify non-Markovian contributions to the noise, all of which are condensed in various simple filtered integrals that, to the best of our knowledge, have not been previously discussed in the literature. We have also provided a compact Pauli-twirled approximation of the channel, which maintains the non-Markovian effect during each driving within a single gate, but can be used in large-scale simulations of noisy Clifford circuits.
@@
We have gauged our analytical approach extensively against direct numerical simulations of the noisy dynamics as well as against map reconstructions based on various approximate models. Comparing the average channel infidelity of single-qubit gates subject to an extended phase diffusion model of the laser, we have found that our non-Clifford and non-Markovian dynamical map outperforms the literature-standard depolarizing noise model. We further see that coherent over-rotations present in the noise can be explained when including non-Markovian contributions to the noise that have previously not been described. Furthermore, we have extended our formalism to handle universal, multi-axis colored noise. This allows a compact methodology to study analytically the effect of multi-axis stochastic fluctuations in the dynamics of single-qubit gates from the characterization of the noise properties alone. Given the broad assumptions of our treatment we have also used this model’s predictions to compare against experimental data obtained via quantum process tomography and randomized benchmarking. For short times, that is the regime of few Rabi flops in which QIPs typically operate, we see that non-Markovian contributions arise, although they are small in comparison to other contributions to the error. However, we point out that it is expected that non-Markovian contributions will become important in longer sequences of gates as are encountered in algorithms or QCVV methods that employ signal amplification. Capturing these contributions will thus be important in understanding the cumulative effect of noise correlations in QIPs for QCVV and applications.
@@
In our trapped-ion experiments, we find that our reconstructed noise model provides a good lower bound for established infidelity estimators such as randomized benchmarking. In particular, we find that our model outperforms previous gate infidelity metrics based on filtered power spectral densities, and captures the main contributions of experimental infidelity. We find as expected that the effect of amplitude fluctuations is subdominant under the assumption that there is no (or little) cross-correlation between phase and amplitude. We note that the resource overhead of quantum process tomography reconstructions and estimation of gate infidelity is substantial for trapped ions, such that they were entirely dominated by projection noise in our study. However, in experimental platforms where experimental shots are more readily accumulated, such as in superconducting qubits, this should remain a viable approach. Our predictions are compared against (standard) randomized benchmarking, which can be inaccurate under several circumstances that are experimentally relevant. In the future, other techniques such as cycle benchmarking [130] could be employed for a more stringent test.
@@ Acknowledgements.
A.B. warmly thanks P. Zoller for sharing the lecture notes in [82], for interesting discussions regarding the fast-fluctuation expansion in multiplicative SDEs including the connection to relevant previous works for the dressed-state formalism in SDEs [89], as well as noting the existence of exact solutions for the driven qubit under a Ornstein-Uhlenbeck frequency noise. A.S. would like to thank M. Meth for his assistance with operating the machine to take randomized benchmarking and QPT data. The project leading to this publication has received funding from the US Army Research Office through Grant No. W911NF-21-1-0007. A.B acknowledges support from PID2021-127726NB- I00 (MCIU/AEI/FEDER, UE), from the Grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033, from the CSIC Research Platform on Quantum Technologies PTI-001, and from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101114305 (“MILLENION-SGA1” EU Project). R.F., A.S., M.G.B., Ch.D.M., and T.M. also acknowledge funding by the IQI GmbH. This research was funded in part by the Austrian Science Fund (FWF) [10.55776/F71]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.